OpenStudy (anonymous):

The circumference of the Earth is approximately 25,000 miles. Find the diameter of the Earth

OpenStudy (apoorvk):

cirumference= 25000 = \[\pi\] times diameter. \[\pi\]=3.14 calculate!!!

OpenStudy (anonymous):

A. 7,957.8 miles B. 7,958.7 miles C. 7,958.6 miles D. 7,957.7 miles

OpenStudy (apoorvk):

okay u mean approximations... hmm wait

OpenStudy (paxpolaris):

are u allowed calculator....?

OpenStudy (anonymous):

Yes

OpenStudy (paxpolaris):

\[C=\pi \cdot D\]\[\large D= \frac C \pi = {25000 \over \pi}\] http://lmgtfy.com/?q=25000%2Fpi so, D. 7,957.7 miles